3. While traveling on the turnpike, a 2004 hybrid car traveled 260 miles on 3.5 gallons of gas. How much gas will be used to travel 350 miles on the turnpike?
Answer by stanbon(75887)

While traveling on the turnpike, a 2004 hybrid car traveled 260 miles on 3.5 gallons of gas. How much gas will be used to travel 350 miles on the turnpike?
-----
Use a proportion::
gas/dist = gas/dist
----------------
x/350 = 3.5/260
-----
x = 350*3.5/260 = 4.71 gallons (gas needed)
===============
Cheers,
Stan H.
================